{2,3,5,7,11,13,17,19,23,29,31,37

Show that Goldbach's Conjecture is true for the even numbers from 20 to 40 by writing each number as a sum of two primes.

Answer by jim_thompson5910(35256)

Here are all the possible sums for each even number from 20 to 40:

3+17=20
7+13=20
17+3=20
3+19=22
5+17=22
11+11=22
5+19=24
7+17=24
11+13=24
3+23=26
7+19=26
13+13=26
5+23=28
11+17=28
7+23=30
11+19=30
13+17=30
3+29=32
13+19=32
3+31=34
5+29=34
11+23=34
17+17=34
5+31=36
7+29=36
13+23=36
17+19=36
7+31=38
19+19=38
3+37=40
11+29=40
17+23=40